# Conserved quantities of distinguished curves on conformal sphere

### v1

@@ 1. Introduction
Our first generalization follows [9], where authors use tractor calculus to construct conserved quantities of conformal circles. Explicitly, they construct a series of canonical tractors defined for each curve and show that the wedge of the first three of them is parallel along the curve only if the curve is a conformal circle. Then pairing with suitable BGG solutions gives conserved quantities of conformal circles. We focus in Section 2 on curves such that the wedge of the first four canonical tractors is parallel. On the conformal round sphere, it is then a direct generalization to find conserved quantities for such curves by pairing with suitable parallel tractors.
@@ 2.2. Curves and tractors
We consider arbitrary curve $\gamma=(\gamma^{i}(t))$, $i=1,\dots,n$, and we denote
@@
The best known and studied class of conformal curves are conformal circles, [8, 9, 14]. In the language of tractors, the curve is a conformal circle if and only if tractors $\mathbb{T},\mathbb{U},\mathbb{A},\mathbb{A^{\prime}}$ are linearly dependent, i.e. $\mathbb{A^{\prime}}=\langle\mathbb{T},\mathbb{U},\mathbb{A}\rangle$. This leads to the condition that $T_{3}=\mathbb{T}\wedge\mathbb{U}\wedge\mathbb{A}$ is parallel along $\gamma$ for the tractor connection, [9]. For each (parametrized) curve, thus for each sequence of tractors $\mathbb{T},\mathbb{U},\mathbb{A},\mathbb{A^{\prime}},\dots$ of arbitrary length $\ell$, we can compute the following Gramm matrix of order $\ell$
@@
The above observations show that the conditions $\Delta_{4}\neq 0$ and $\Delta_{5}=0$ are equivalent to the linear dependency of the tractors $\mathbb{T},\mathbb{U},\mathbb{A},\mathbb{A^{\prime}}$ and $\mathbb{A^{\prime\prime}}$ and the explicit combination takes form (2.4). Note that $\Delta_{4}$ is non-positive, [14, Lemma 2.3.]. The computation (LABEL:derivative-4-tractor) then gives that (2.6) is parallel with respect to the tractor connection. Conversely, assuming (2.6) is parallel, we get
@@
Since $\Delta_{4}$ is a relative invariant, the condition that $\Delta_{4}$ is non-zero constant gives a special class of (parametrized) curves. In particular, we can normalize the constant to $\Delta_{4}=-1$, which corresponds to the conformal arc-length parametrization.
@@ 2.3. Conserved quantities and tractors
In [9], authors give a general method, how to use conformal Killing-Yano $2$-forms and the parallel $3$-tractor $\mathbb{T}\wedge\mathbb{U}\wedge\mathbb{A}$ to find conserved quantities of conformal circles on arbitrary conformal manifolds. The situation on $S^{n}$ is simpler; let us outline here the situation. For interest, we give the resulting quantities on $S^{n}$ in Appendix A.
@@
$Q^{ijkp}$ $=u^{-4}\epsilon^{ijkp}(X,U,A,A^{\prime})$ (2.12)
@@
In the coordinates $w^{a}$, we can write the corresponding transformation matrix $\rho_{4}(J)$ and thus $\exp(-\rho_{4}(J))$
$\rho_{4}(J)=\left(\begin{smallmatrix}0&0&0&0&0\\ -X^{3}&0&0&0&0\\ X^{2}&0&0&0&0\\ -X^{1}&0&0&0&0\\ 0&X^{3}&-X^{2}&X^{1}&0\end{smallmatrix}\right),\ \ \exp(-\rho_{4}(J))=\left(\begin{smallmatrix}1&0&0&0&0\\ -X^{3}&1&0&0&0\\ X^{2}&0&1&0&0\\ -X^{1}&0&0&1&0\\ -\frac{1}{2}\big{(}(X^{1})^{2}+(X^{2})^{2}+(X^{3})^{2}\big{)}&X^{3}&-X^{2}&X^{1}&1\end{smallmatrix}\right).$
@@
$Q^{0124}=u^{-4}X^{3}\cdot U\wedge A\wedge A^{\prime}-u^{-2}\>U\wedge A^{\prime}+3u^{-4}\>\langle U,A\rangle\>U\wedge A$
$Q^{0134}=u^{-4}X^{2}\cdot U\wedge A\wedge A^{\prime}-u^{-2}\>U\wedge A^{\prime}+3u^{-4}\>\langle U,A\rangle\>U\wedge A$
$Q^{0234}=u^{-4}X^{1}\cdot U\wedge A\wedge A^{\prime}-u^{-2}\>U\wedge A^{\prime}+3u^{-4}\>\langle U,A\rangle\>U\wedge A$
@@
Let us note that one of the possible definitions of logarithmic spirals is via invariants, [14]. We say that the curve is a spiral if $\kappa_{1}$ is constant and $\Delta_{5}=0$, where
@@ 3.1. Conformal Mercator equation in the conformally-flat case
Choosing the Euclidean metric $\delta$ from the conformal class, the conformal Mercator equation [5] is given by
@@ 3.2. Conserved quantities from conformal Killing vector fields
Based on Noether’s principle, let us now derive an expression for the conserved quantities for solutions of the conformal Mercator equation (3.15). Let $\gamma$ be a solution of (3.15) and $\{\phi_{s}\}$ a one-parameter group of symmetries of the Lagrangian (3.16). Consider a one-parameter family of solution curves $\Gamma:[s_{0},s_{1}]\times[t_{0},t_{1}]\to\mathbb{R}^{n}$ given by $\Gamma(s,t)=\phi_{s}(\gamma(t)).$ Choose the parameter $s$ such that $0\in[s_{0},s_{1}]$ and $\phi_{0}={\rm Id},$ i.e. $\Gamma(0,t)=\gamma(t).$ If we restrict the range of the parameter $t$ so that the generator of $\{\phi_{s}\}$ is nowhere tangent to $\gamma,$ then for a sufficiently small range of $s$ around $0,$ the family $\Gamma$ is a $2$-dimensional submanifold of $\mathbb{R}^{n}.$
@@
and $C$ is defined in (3.15). Let us note that in obtaining (3.18), we have used the flatness of $\delta$ and the fact that $[U,V]=0.$ These imply that $\nabla_{V}U=\nabla_{U}V$ and $\nabla_{V}\nabla_{U}U=\nabla_{U}\nabla_{V}U.$
@@
The proposition below follows from the fact that the Lagrangian (3.16) is invariant under a one-parameter group of conformal transformations. [^4] One can verify directly that the Lagrangian (3.16) satisfies (3.17) for any one-parameter group conformal transformations $\phi_{s}.$
@@ 3.3. Expressions for conserved quantities and examples
${\mathcal{Y}}$ $=$ $u^{-2}\langle U,X\rangle A-\left(1+u^{-2}\langle A,X\rangle\right)U\left(u^{-2}\langle U,A\rangle+\langle C,X\rangle\right)X-\frac{|X|^{2}}{2}C,$
@@ 4.1. Hamiltonian formalism
The Hamiltonian is then given by the Legendre transform ${H=\langle\mathcal{P},X\rangle+\langle\mathcal{R},U\rangle-L_{1},}$ which gives
@@ Appendix A Conserved quantities for conformal circles from parallel $3$-tractor
Let us give here for completeness the quantities for conformal circles coming from pairing of the $3$-tractor $\mathbb{T}\wedge\mathbb{U}\wedge\mathbb{A}$ and parallel tractors corresponding to CKY 2-forms aka elements of $\wedge^{3}\mathbb{R}^{n+2}$. Computations analogous to Section 2.3 give that the $3$-tractor takes form
@@ Appendix C Note on tractor equation
The general principle states, that the middle aka invariant slot of the tractor equation
@@
It turns out that involving the restriction $\alpha^{\prime}_{1}=0$ on this expression results in the minus $u^{-1}$-multiple of (LABEL:eq). This agrees with the fact, that the tractor equation lives in the space $TM[-1]$ while the flat Mercator equation lives in the space $TM[-2]$. Let us note that this works only in the case of a conformal sphere. In the general curved case, the Mercator equation differs from the equation coming from tractor calculations by some curvature term.

### v2

@@ 1. Introduction
Our first generalization follows [9], where authors used tractor calculus to construct conserved quantities of conformal circles. Explicitly, they constructed a series of canonical tractors defined for each curve and showed that the wedge of the first three of them is parallel along the curve only if the curve is a conformal circle. Then pairing with suitable BGG solutions gives conserved quantities of conformal circles. We focus in Section 2 on curves such that the wedge of the first four canonical tractors is parallel. On the conformal round sphere, it is then a direct generalization to find conserved quantities for such curves by pairing with suitable parallel tractors.
@@ 2.2. Curves and tractors
We consider an arbitrary curve $\gamma=(\gamma^{i}(t))$, $i=1,\dots,n$, and we denote
@@
The best known and studied class of conformal curves are conformal circles, [8, 9, 14]. In the language of tractors, a curve is a conformal circle if and only if tractors $\mathbb{T},\mathbb{U},\mathbb{A},\mathbb{A^{\prime}}$ are linearly dependent, i.e. $\mathbb{A^{\prime}}=\langle\mathbb{T},\mathbb{U},\mathbb{A}\rangle$. This leads to the condition that $T_{3}=\mathbb{T}\wedge\mathbb{U}\wedge\mathbb{A}$ is parallel along $\gamma$ for the tractor connection, [9]. For each (parametrized) curve, thus for each sequence of tractors $\mathbb{T},\mathbb{U},\mathbb{A},\mathbb{A^{\prime}},\dots$ of arbitrary length $\ell$, we can compute the following Gramm matrix of order $\ell$
@@
The above observations show that the conditions $\Delta_{4}\neq 0$ and $\Delta_{5}=0$ are equivalent to the linear dependency of the tractors $\mathbb{T},\mathbb{U},\mathbb{A},\mathbb{A^{\prime}}$ and $\mathbb{A^{\prime\prime}}$ and the explicit combination takes the form (2.4). Note that $\Delta_{4}$ is non-positive, [14, Lemma 2.3.]. The computation (LABEL:derivative-4-tractor) then gives that (2.6) is parallel with respect to the tractor connection. Conversely, assuming (2.6) is parallel, we get
@@
Since $\Delta_{4}$ is a relative invariant, the condition that $\Delta_{4}$ is a non-zero constant gives a special class of (parametrized) curves. In particular, we can normalize the constant to $\Delta_{4}=-1$, which corresponds to the conformal arc-length parametrization.
@@ 2.3. Conserved quantities and tractors
In [9], authors give a general method of how to use conformal Killing-Yano $2$-forms and the parallel $3$-tractor $\mathbb{T}\wedge\mathbb{U}\wedge\mathbb{A}$ to find conserved quantities of conformal circles on arbitrary conformal manifolds. The situation on $S^{n}$ is simpler; let us outline here the situation. For interest, we give the resulting quantities on $S^{n}$ in Appendix A.
@@
$Q^{ijkl}$ $=u^{-4}\epsilon^{ijkl}(X,U,A,A^{\prime})$ (2.12)
@@
In the coordinates $w^{a}$, we can write the corresponding transformation matrix $\rho_{4}(J)$ and thus $\exp(-\rho_{4}(J))$ as
$\rho_{4}(J)=\left(\begin{smallmatrix}0&0&0&0&0\\ -X^{3}&0&0&0&0\\ X^{2}&0&0&0&0\\ -X^{1}&0&0&0&0\\ 0&X^{3}&-X^{2}&X^{1}&0\end{smallmatrix}\right),\ \ \exp(-\rho_{4}(J))=\left(\begin{smallmatrix}1&0&0&0&0\\ X^{3}&1&0&0&0\\ -X^{2}&0&1&0&0\\ X^{1}&0&0&1&0\\ -\frac{1}{2}\big{(}(X^{1})^{2}+(X^{2})^{2}+(X^{3})^{2}\big{)}&-X^{3}&X^{2}&-X^{1}&1\end{smallmatrix}\right).$
@@
$Q^{0124}=u^{-4}X^{3}\,U\wedge A\wedge A^{\prime}-u^{-2}\>U\wedge A^{\prime}+3u^{-4}\>\langle U,A\rangle\>U\wedge A$
$Q^{0134}=u^{-4}X^{2}\,U\wedge A\wedge A^{\prime}-u^{-2}\>U\wedge A^{\prime}+3u^{-4}\>\langle U,A\rangle\>U\wedge A$
$Q^{0234}=u^{-4}X^{1}\,U\wedge A\wedge A^{\prime}-u^{-2}\>U\wedge A^{\prime}+3u^{-4}\>\langle U,A\rangle\>U\wedge A$
@@
Let us note that one of the possible definitions of logarithmic spirals is via invariants, [14]. We say that a curve is a spiral if $\kappa_{1}$ is constant and $\Delta_{5}=0$, where
@@ 3.1. Conformal Mercator equation in the conformally-flat case
Choosing the Euclidean metric $\delta$ in the conformal class, the conformal Mercator equation [5] is given by
@@ 3.2. Conserved quantities from conformal Killing vector fields
Based on Noether’s principle, let us now derive an expression for the conserved quantities of solutions of the conformal Mercator equation (3.15). Let $\gamma$ be a solution of (3.15) and $\{\phi_{s}\}$ a one-parameter group of symmetries of the Lagrangian (3.16). Consider a one-parameter family of solution curves $\Gamma:[s_{0},s_{1}]\times[t_{0},t_{1}]\to\mathbb{R}^{n}$ given by $\Gamma(s,t)=\phi_{s}(\gamma(t)).$ Choose the parameter $s$ such that $0\in[s_{0},s_{1}]$ and $\phi_{0}={\rm Id},$ i.e. $\Gamma(0,t)=\gamma(t).$ If we restrict the range of the parameter $t$ so that the generator of $\{\phi_{s}\}$ is nowhere tangent to $\gamma,$ then for a sufficiently small range of $s$ around $0,$ the family $\Gamma$ is a $2$-dimensional submanifold of $\mathbb{R}^{n}.$
@@
and the expression for $C$ is given by (3.15). Let us note that in obtaining (3.18), we have used the flatness of $\delta$ and the fact that $[U,V]=0.$ These imply that $\nabla_{V}U=\nabla_{U}V$ and $\nabla_{V}\nabla_{U}U=\nabla_{U}\nabla_{V}U.$
@@
The proposition below follows from the fact that the Lagrangian (3.16) is invariant under a one-parameter group of conformal transformations. [^4] One can verify directly that the Lagrangian (3.16) satisfies (3.17) for any one-parameter group of conformal transformations.
@@ 3.3. Expressions for conserved quantities and examples
${\mathcal{Y}}$ $=$ $u^{-2}\langle U,X\rangle A-\left(1+u^{-2}\langle A,X\rangle\right)U+\left(u^{-2}\langle U,A\rangle+\langle C,X\rangle\right)X-\frac{|X|^{2}}{2}C,$
@@ 4.1. Hamiltonian formalism
The Hamiltonian is then given by the Legendre transform ${H=\langle\mathcal{P},X^{\prime}\rangle+\langle\mathcal{R},U^{\prime}\rangle-L_{1},}$ which gives
@@ Appendix A Conserved quantities for conformal circles from parallel $3$-tractor
Let us give here for completeness the quantities for conformal circles coming from pairing of the $3$-tractor $\mathbb{T}\wedge\mathbb{U}\wedge\mathbb{A}$ and parallel tractors corresponding to CKY 2-forms aka elements of $\wedge^{3}\mathbb{R}^{n+2}$. Computations analogous to Section 2.3 give that the $3$-tractor takes the form
@@ Appendix C Note on tractor equation
The general principle states that the middle aka invariant slot of the tractor equation
@@
It turns out that involving the restriction $\alpha^{\prime}_{1}=0$ on this expression results in the minus $u^{-1}$-multiple of (LABEL:eq). This agrees with the fact that the tractor equation lives in the space $TM[-1]$ while the flat Mercator equation lives in the space $TM[-2]$. Let us note that this works only in the case of the conformal sphere. In the general curved case, the Mercator equation differs from the equation coming from tractor calculations by some curvature term.